OpenStudy (anonymous):

Find the domain of g[g(X)] if f(x)= x^2+4 and g(x) = sqrt{x+2} Find the domain of g[g(X)] if f(x)= x^2+4 and g(x) = sqrt{x+2} @Mathematics

OpenStudy (anonymous):

\[g(g(x))=g(\sqrt{x+2})=\sqrt{\sqrt{x+2}+2}\]

OpenStudy (anonymous):

oh you can use a double radical?

OpenStudy (anonymous):

domain you find by solving \[x+2\geq 0\] so \[x\geq -2\]

OpenStudy (anonymous):

awesome, you are smart! thanks